OpenStudy (king):

Help!!...........Let a,b,c be real numbers such that a-7b+8c=4 and 8a+4b-c=7.Find a^2-b^2+c^2 Help!!...........Let a,b,c be real numbers such that a-7b+8c=4 and 8a+4b-c=7.Find a^2-b^2+c^2 @Mathematics

OpenStudy (king):

Its 8a+4b-c=7

OpenStudy (asnaseer):

\[a^{2}-b^{2}+c^{2}=0\]

OpenStudy (asnaseer):

It's a fairly lengthy proof

OpenStudy (asnaseer):

I made a mistake somewhere - ignore the above These are the steps I have so far: a-7b+8c=4 8a+4b-c=7 using elimination you can get the following 3 equations in just 2 variables each: (1) 13c-12b=5 (2) 13a+5b=12 (3) 12a+5c=13 (1)*(2), (1)*(3), and (2)*(3) then gives: (4) 13*13ac+5*13bc-12*13ab-5*12b^2=5*12 (5) 12*13ac+5*13c^2-12*12ab-5*12bc=5*13 (6) 12*13a^2+5*13ac+5*12ab+5*5bc=12*13 then 5*(6)+13*(4)+12*(5) gives: 5*12*13(a^2-b^2+c^2) +(5*5*13+13*13*13+12*12*13)ac+(5*5*12-12*13*13-12*12*12)ab+(5*5*5+5*12*13-5*12*12)bc=5*12.13+5*12*13+5*12*13=3*5*12*13 this ca be re-written as: 5*12*13(a^2-b^2+c^2)+13(5*5+13*13+12*12)ac+12(5*5-13*13-12*12)ab+5(5*5+12*13-12*12)bc=3*5*12*13 giving: a^2-b^2+c^2 = (3*5*12*13 - 5(5*5+12*13-12*12)bc - 12(5*5-13*13-12*12)ab - 13(5*5+13*13+12*12)ac) / 5*12*13

OpenStudy (asnaseer):

\[a^{2}-b^{2}+c^{2}=3-\frac{5(5^{2}-12^{2}+13^{2})bc+12(5^{2}-12^{2}-13^{2})ab+13(5^{2}+12^{2}+13^{2})ac}{5*12*13}\]

OpenStudy (asnaseer):

I'm sure I still have some mistakes - tricky question! :-)

OpenStudy (zarkon):

I get a^2-b^2+c^2=1

OpenStudy (zarkon):

solve a-7b+8c=4 and 8a+4b-c=7 you get \[a=\frac{13}{12}-\frac{5}{12}t\] \[b=\frac{-5}{12}+\frac{13}{12}t\] \[c=t\] then compute \[a^2-b^2+c^2\] you will get 1

OpenStudy (king):

I still dont understand do you mind explaining without wrting those \frac.....

OpenStudy (zarkon):

I solved the system a-7b+8c=4 and 8a+4b-c=7 ( I row reduced the augmented matrix and solved in terms of the free variable)

OpenStudy (king):

Can you solve this sum without matrices as i have not learnt that topic yet...... :(

OpenStudy (zarkon):

anything you can do with matrices you can do without (it just usually takes longer)

OpenStudy (king):

So,can you tell me how to do it without it

OpenStudy (zarkon):

a-7b+8c=4 a=7b-8c+4 8a+4b-c=7 8(7b-8c+4)+4b-c=7 solve for b b=(13c-5)/12 replace b in a=7b-8c+4 with b=(13c-5)/12 a=7[(13c-5)/12]-8c+4 =(13-5c)/12 now compute \[a^2-b^2+c^2\] \[\left[(13-5c)/12\right]^2-\left[(13c-5)/12\right]^2+c^2\]